OLIVER69:

Suppose an experiment has 5 outcomes. Outcome 1 is twice as likely as outcome 2. Outcome 2 is twice as likely as outcome 3. Outcome 3 is twice as likely as outcome 4. Outcome 4 is twice as likely as outcome 5. Find the probability of outcome 1. [Write your answer as a fraction. For example, 1/2 or 2/3 or 15/16, etc.] [Hint: Call outcome 5 the "x" and write all other outcomes in terms of "x" to find the solution.]

Markthegoat1117:

Let's assign a variable to the probability of outcome 5, which is represented by "x". According to the given information, we can express the probabilities of the other outcomes in terms of "x": Outcome 4: The probability of outcome 4 is twice as likely as outcome 5, so it is 2x. Outcome 3: The probability of outcome 3 is twice as likely as outcome 4, so it is 2(2x) = 4x. Outcome 2: The probability of outcome 2 is twice as likely as outcome 3, so it is 2(4x) = 8x. Outcome 1: The probability of outcome 1 is twice as likely as outcome 2, so it is 2(8x) = 16x. To find the probability of outcome 1, we need to sum up the probabilities of all outcomes, which should equal 1: x + 2x + 4x + 8x + 16x = 1 Simplifying the equation: 31x = 1 Dividing both sides by 31: x = 1/31 Therefore, the probability of outcome 1 is 16x, which is 16 * (1/31) = 16/31.

OLIVER69:

Thanks so much for this wonderful explanation.

Markthegoat1117:

@oliver69 wrote:
Thanks so much for this wonderful explanation.
Your welcome happy to help😁